Write the slope-intercept equation for the line with slope 4/5 and y-intercept (0, -1) and graph the line. Submit your graph to the Dropbox. Please show all of your work.
Answer by mananth(16949)

m= 4/ 5

Plug value of the slope and point ( 0 , -1 ) in
y=mx+b
-1.00 = 4/5 * 0 +b
b= -1 + -0

b= -1
So the equation will be
Y = 4/5 x -1
Assume values of x and plug in the equation to get values of y.
Plot the points
x 0 5 -5
y= -1 3 -5
(x,y) (0,-1),(5,3),( -5,-5)




m.ananth@hotmail.ca